write the expression for the instantaneous rate of the following reaction in term of changes in concentration of O3 or O2 : 2 O3 (g) - 3 O2 (g)
1
Expert's answer
2019-09-11T11:40:00-0400

"2 O_3 (g) \to 3 O_2 (g)"

Instantaneous reaction rate in terms of change in concentration:

"-\frac{1}{2}\frac{d[O_3]}{dt}=\frac{1}{3}\frac{d[O_2]}{dt}"

Negative sign indicates the decreasing concentration of ozone gas.